# Dark Photon Dark Matter and Low-Frequency Gravitational Wave Detection with Gaia-like Astrometry

### v1

@@ Abstract
Astrometric surveys offer us a method to search for elusive cosmic signatures, such as ultralight dark photon dark matter and gravitational waves, by observing the deflection to the apparent positions of the stars. The detection capabilities of such surveys rapidly decrease at low frequencies, because the signals become hardly distinguishable from the background motion of stars. In this work, we find that the background motion can be well described by a linear model over time, based on which we propose a linear background subtraction scheme. Compared to the conventional quadratic subtraction, the advantage of linear subtraction emerges within the frequency range below $6\times 10^{-9}~{}{\rm Hz}$. Taking dark photons with purely gravitational interactions, dark photons with additional $U(1)_{B}$ or $U(1)_{B-L}$ gauge interactions, and low-frequency gravitational waves as examples, we illustrate that the linear subtraction scheme can result in an enhancement of more than one order of magnitude in the exclusion limits of Gaia-like experiments in the low-frequency range.
@@ I Introduction
The development of astrophysical observation methods has revolutionized our understanding of the cosmos, enabling us to probe subtle effects induced by new astrophysical sources or new physics with unprecedented precision. Recently, several Pulsar Timing Array (PTA) collaborations reported first evidence for the stochastic gravitational wave (GW) background at nano-Hertz (nHz) frequencies  Allen (1996); Agazie et al. (2023a); Antoniadis et al. (2023); Reardon et al. (2023); Xu et al. (2023), which may arise from a population of inspiraling supermassive black hole binaries Sesana et al. (2008).
@@
Apart from the GW signals, PTA experiments can also probe several new physics models. One example would be the ultralight bosonic dark matter Schive et al. (2014); Hui et al. (2017) which is capable of solving the so called small-scale problems Weinberg et al. (2015) of the cold dark matter model. Ultralight dark photon dark matter (DPDM) assumes that dark matter is composed of vector fields Adshead and Lozanov (2021); Amin et al. (2022) with a mass around $10^{-22}~{}{\rm eV}$. Due to its wave nature, the field value of the ultralight DPDM oscillates in an approximate coherent manner. Therefore, the ultralight DPDM can induce nontrivial oscillations of the metric due to the pure gravitational effects Khmelnitsky and Rubakov (2014); Nomura et al. (2020); Chen et al. (2023), or impose external oscillating forces on the test objects via additional U(1) gauge interactions Pierce et al. (2018), both of which can modify the pulse arrival time Porayko et al. (2018); Xue et al. (2022); Afzal et al. (2023); Nelson and Scholtz (2011); Wu et al. (2022); Xia et al. (2023).
@@
On the other hand, the GW or DPDM background can result in the deflection of light emitted from the astrophysical objects to the Earth, due to the following two distinct effects:
@@
  • GWs or DPDM with purely gravitational interactions can induce metric perturbations, which can affect the trajectory of light emitted from the sources Braginsky et al. (1990); Pyne et al. (1996); Kaiser and Jaffe (1997); Kopeikin et al. (1999); Book and Flanagan (2011). As a result, the apparent positions of stars we observe slightly deviate from their initial positions. Such displacements depend on metric perturbations at both the Earth and stars, while the Earth term should be dominant, as we will explain later. It should also be mentioned that, unlike the deflection induced by ultralight scalar dark matter, which is suppressed due to an approximate spherical symmetry, the deflection caused by DPDM is generally not suppressed.
@@
  • DPDM coupling to the baryon number ${\rm U}(1)_{B}$ or the baryon number minus lepton number ${\rm U}(1)_{B-L}$ can lead to the aberration effect, which also changes the apparent position of stars Guo et al. (2019).
@@
By precisely measuring the variation of stars’ apparent positions, it is hopeful to capture the traces of GW and DPDM signals in the Gaia mission Prusti et al. (2016) and upcoming astrometric surveys like Roman Wang et al. (2021, 2022); Haiman et al. (2023); Pardo et al. (2023) and Theia Boehm et al. (2017); Malbet et al. (2021); Garcia-Bellido et al. (2021). Gaia observes positions, proper motions, and parallaxes of more than $10^{9}$ objects with unprecedented accuracy Prusti et al. (2016). The sensitive frequency of Gaia is determined by its survey lifetime $T_{\rm obs}=10~{}{\rm year}$ and observational cadence $\Delta t=24~{}{\rm day}$, typically, $10^{-9}~{}{\rm Hz}\lesssim f\lesssim 10^{-7}~{}{\rm Hz}$. Hence Gaia can provide complementary detection capabilities to PTA experiments for nHz astrophysical sources Moore et al. (2017); Klioner (2018); Qin et al. (2019); O’Beirne and Cornish (2018); Mihaylov et al. (2018); Bini and Geralico (2018); Darling et al. (2018); Mihaylov et al. (2020); Guo et al. (2019); Qin et al. (2021); Aoyama et al. (2021); Jaraba et al. (2023); Liang et al. (2024); Çalışkan et al. (2024).
@@
To extract the deflections to the apparent positions of stars caused by external signals, it is crucial to subtract their intrinsic background motion stemming from the physical motion of stars with respect to the solar barycenter and the secular aberration due to the moving reference system. This intrinsic motion can induce large angular deflection far exceeding the resolution of Gaia. For the signals with frequencies $f\gtrsim 2/T_{\rm obs}\sim 6\times 10^{-9}~{}{\rm Hz}$, the background subtraction essentially does not influence the sensitivity, since the oscillation pattern of GWs or DPDM is distinctive from the background. Conversely, if $f\lesssim 6\times 10^{-9}~{}{\rm Hz}$, a half oscillating cycle would not be observable in the time-series data within the operational lifetime of Gaia. Then the signal is also partially subtracted, thereby reducing the sensitivity of detecting low-frequency GW and DPDM signals. Analogous to PTAs, the quadratic model is commonly used to subtract the background proper motion in the existing literature Moore et al. (2017); Guo et al. (2019); Wang et al. (2021, 2022). However, a more comprehensive study of the background motion of stars and subtraction schemes should be implemented.
@@
In this work, we propose a linear background subtraction scheme for the time-series data of apparent stellar positions in Gaia-like astrometric surveys. Upon analyzing the background motion of stars, we find that the impact on the apparent stellar position due to the constant acceleration from the galactic gravitational potential is negligibly small relative to Gaia’s sensitivity. Meanwhile, the galactocentric acceleration of the solar barycenter mimics a linear term over time in the proper motion due to aberration. These observations serve as the primary basis for our linear subtraction approach. Nevertheless, unresolvable binary systems could bring additional noise, which cannot be effectively subtracted. We provide concrete examples by calculating the astrometric deflection induced by DPDM with purely gravitational interactions, while also revisiting scenarios involving DPDM with ${\rm U}(1)_{B}$ or ${\rm U}(1)_{B-L}$ gauge symmetry and GWs from binary systems. Subsequent numerical simulation using mock data reveals that our linear subtraction scheme significantly enhances Gaia’s detection capabilities in the low-frequency range, improving sensitivity by more than an order of magnitude compared to the quadratic subtraction.
@@
The layout of this paper is as follows. In Sec. II we illustrate the background motion of stars can be well described by a linear model over time. In Sec. III we calculate astrometric deflections induced by GWs and DPDM. The numerical simulation and results are presented in Sec. IV. We summarize our main conclusions in Sec. V.
II Background motion of the apparent stellar locations
@@
Astrometric objects have background motion around the galactic center. Their motion is randomly distributed with a potentially non-zero expectation value depending on the star’s sky location, due to the common motion of the galactic disk. We assume that the velocity $\boldsymbol{v}=(v_{r},v_{\theta},v_{\phi})$ in spherical polar coordinates at the galactic center follows a Gaussian distribution
@@
where $\boldsymbol{n}$ denotes the sky location of stars, $v_{0,i}(\boldsymbol{n})$ represents the expectation value of $\boldsymbol{v}$ in each direction, and $\sigma_{v,i}$ refers to the velocity dispersion. The anisotropy of the velocity dispersion has been observed by SDSS and Gaia. Approximately we have the relation $\sigma_{v,r}\simeq 2\sigma_{v,\theta/\phi}$ for stars with high metallicity, which may originates from a major merger event of a satellite galaxy between $8$ and $11$ Gyr ago Belokurov et al. (2018). For qualitative discussions, we assume the uncertainty $\sigma_{v}\simeq 100\,{\rm km}/{\rm s}\simeq 1.01\times 10^{-4}\,{\rm pc/year}$ for all stars on both directions on the sky sphere. Choosing a typical distance of $1$ kpc, [^1] The peak distribution of distances between the stars and us observed by Gaia is $1\sim 2$ kpc Vallenari et al. (2023). the angular deflection observed over $10$ years reaches $\simeq 1.01\times 10^{-6}\,{\rm rad}$, which significantly exceeds the single-exposure resolution $\sigma_{\rm ss}\simeq 100\,\mu{\rm as}=4.85\times 10^{-10}\,{\rm rad}$ of Gaia for a single star. It then becomes necessary to subtract the background proper motion from the data before searching for the GW or DPDM signal.
Let us consider stars in two categories, isolated stars and the ones in binary or higher-order systems. For isolated stars, their velocities in general change over time. Hence we should also estimate their accelerations, which mainly include contributions from two aspects:
  • The constant accelerations of stars. The density gradient in the Milky Way could induce constant accelerations for stars. Unlike their proper motion, the constant accelerations of stars are predictable due to the galactic mass profile, indicating the uncertainty $\sigma_{a}$ should be much smaller than the expectation values $|\boldsymbol{a}_{0}|$. Typical stars in the Milky Way experience an acceleration of $|\boldsymbol{a}_{0}|\sim 10^{-10}\,{\rm m}/{\rm s}^{2}$ Silverwood and Easther (2019), which leads to a displacement of $|\boldsymbol{a}_{0}|\Delta t^{2}/2\simeq 1.6\times 10^{-10}$ pc over the 10-year observation. For a star located 1 kpc away from us, the corresponding angular deflection is about $0.03\,\mu{\rm as}$, significantly below the detection threshold.
  • The constant acceleration of solar system barycenter. Like other stellar systems, the velocity of our solar system is not precisely constant. Instead, it follows a curved orbit in the Galaxy. For the observer, the nonzero acceleration of the solar system barycenter gives rise to the aberration effect, which can be considered as extra (spurious) proper motion superimposed on stars Kopeikin and Makarov (2006); Klioner et al. (2021)
@@
    where $\delta\boldsymbol{n}$ denotes the displacement in the apparent stellar position due to aberration, and $\boldsymbol{a}_{\rm SSB}$ is the acceleration of the solar system barycenter (SSB). It then becomes evident that $\boldsymbol{a}_{B}$ only introduces a modification that is linearly dependent on time.
In short, the background motion of the apparent stellar locations is governed by the proper motion of stars, which can be significantly greater than the resolution of Gaia. The acceleration of stars is very minor and thus plays a negligible role. Moreover, the acceleration of the solar system barycenter yields spurious proper motion. The above backgrounds can be eliminated through a linear subtraction scheme.
On the other hand, the unresolved binary or higher-order systems in the Galaxy could produce a significant background motion for the relevant stars. In the practical numerical calculation, we employ a data compression technique akin to the one described in Refs. Moore et al. (2017); Klioner (2018), since the vast number of objects ($\sim 10^{9}$) in the full dataset of Gaia severely slows down the process of numerical simulation. Specifically, we compress the data of $10^{5}$ stars in the same sky area into one virtual star. For each star, the probability of being part of a binary or high-order system and having specific orbital parameters is statistically independent of the status of the other stars. After data compression, the astrometric data of $10^{5}$ stars in each sky area are added together. The background motion of individual stars in binary or higher-order systems manifests as a time-dependent noise of the compressed virtual star.
A survey of nearby ($<250$ pc) stellar objects finds $\sim 55\%$ solar-type stars are isolated, while the remaining $\sim 45\%$ stars are in binary or higher-order systems Raghavan et al. (2010). Here we take the binary systems for simplicity. Following the Kepler’s law, the orbital period $P_{b}$ has the following relation with the semi-major axis $a_{b}$
@@
where $m_{1}$ labels the mass of the primary star and $q=m_{2}/m_{1}<1$. If both stars are observed by Gaia, the neat contribution from the binary to the displacement of stars will approximate to
$r_{\rm neat}\simeq a_{b}(1+e_{b})\frac{1-q}{1+q}\;,$ (4)
where $e_{b}$ is the orbit eccentricity. Then the neat contribution will be fully canceled if $q=1$.
Figure 1: The result of 10000 times simulation of the injected white noise (WN), proper motion (PM) and binary motion. Dashed lines correspond to the cases with white noise removed. The input Gaia operational parameters align with the discussion in Sec. IV.
For the $q<1$ case, we illustrate by considering a $e_{b}=0$ binary composed of two stars with the total mass $M=m_{1}+m_{2}$. The neat amplitude of the observed angular deflection in one direction is
$\begin{split}A_{b}=\frac{r_{\rm neat}\lambda(\iota_{b},\psi_{b})}{\sqrt{2}D}&=6.10\times 10^{-9}{\rm rad}\left(\frac{1-q}{1+q}\right)\frac{\lambda(\iota_{b},\psi_{b})}{\sqrt{2}}\\ &\times\left[\frac{P_{b}}{\rm year}\right]^{2/3}\left[\frac{D}{1{\rm kpc}}\right]^{-1}\left[\frac{M}{2M_{\odot}}\right]^{1/3}\;,\end{split}$ (5)
where $D$ is the distance between the observer and the binary barycenter. Also, $\lambda(\iota_{b},\psi_{b})=\cos\iota_{b}\sin\psi_{b}+\cos\psi_{b}$ with $\iota_{b}$ and $\psi_{b}$ being the inclination angle and the polarization angle of the stellar binary plane, respectively. Eq. (5) indicates binaries with $P_{b}\gtrsim 1\,{\rm year}$ and $D\sim 1\,{\rm kpc}$ may be resolved by Gaia. According to Raghavan et al. (2010); Simonetti et al. (2020), the period distribution approximately follows the log-Normal distribution
@@
We calculate the power spectral density (PSD) originating from binary stars $S_{\rm bin}(f)$ by implementing a Monte-Carlo simulation, where we assume $45\%$ of stars are in binaries. We use Eq. (5) with $D=1$ kpc, while the orbital period is drawn from Eq. (6). Further, $\cos\iota_{b}$, $\psi_{b}$ and $q$ are treated as free parameters, randomly selected from uniform distributions within the intervals of $[-1,1)$, $[0,2\pi)$ and $[0.1,1)$, respectively. In each simulation, we generate astrometric data of $10^{5}$ stars, which are then compressed into one virtual star. We calculate the PSD of astrometric data of the virtual star. The final $S_{\rm bin}(f)$ is obtained by averaging the PSD over all simulations. The total noise PSD $S_{n}(f)$ is the combination of $S_{\rm bin}(f)$ and that from the instrumental noise $S_{w}(f)$, namely,

### v2

@@ Abstract
Astrometric surveys offer us a method to search for elusive cosmic signatures, such as ultralight dark photon dark matter and gravitational waves, by observing the temporal change of stars’ apparent location. The detection capabilities of such surveys rapidly decrease at low frequencies, because the signals become hardly distinguishable from the background motion of stars. In this work, we find that the background motion can be well described by a linear model over time, based on which we propose a linear background subtraction scheme. Compared to the conventional quadratic subtraction, the advantage of linear subtraction emerges within the frequency range below $6\times 10^{-9}~{}{\rm Hz}$. Taking dark photons with purely gravitational interactions, dark photons with additional $U(1)_{B}$ or $U(1)_{B-L}$ gauge interactions, and low-frequency gravitational waves as examples, we illustrate that the linear subtraction scheme can result in an enhancement of more than one order of magnitude in the exclusion limits of Gaia-like experiments in the low-frequency range.
@@ 1 Introduction
The development of astrophysical observation methods has revolutionized our understanding of the cosmos, enabling us to probe subtle effects induced by new astrophysical sources or new physics with unprecedented precision. Recently, several Pulsar Timing Array (PTA) collaborations reported first evidence for the stochastic gravitational wave (GW) background at nano-Hertz (nHz) frequencies  (Allen (1996); Agazie et al. (2023a); Antoniadis et al. (2023); Reardon et al. (2023); Xu et al. (2023)), which may arise from a population of inspiraling supermassive black hole binaries (Sesana et al. (2008)).
@@
Apart from the GW signals, PTA experiments can also probe several new physics models. One example would be the ultralight bosonic dark matter (Schive et al. (2014); Hui et al. (2017)) which has the potential to solve the so called small-scale problems (Weinberg et al. (2015)) of the cold dark matter model. Ultralight dark photon dark matter (DPDM) assumes that dark matter is composed of vector fields (Adshead & Lozanov (2021); Amin et al. (2022)) with a mass around $10^{-22}~{}{\rm eV}$. Due to its wave nature, the field value of the ultralight DPDM oscillates in an approximately coherent manner. Therefore, the ultralight DPDM can induce nontrivial oscillations of the metric due to the pure gravitational effects (Khmelnitsky & Rubakov (2014); Nomura et al. (2020); Chen et al. (2023); Yu et al. (2024)), or impose external oscillating forces on the test objects via additional U(1) gauge interactions (Pierce et al. (2018)), both of which can modify the pulse arrival time (Porayko et al. (2018); Xue et al. (2022); Afzal et al. (2023); Nelson & Scholtz (2011); Wu et al. (2022); Xia et al. (2023); Chowdhury et al. (2023)).
@@
The GW or DPDM background can also result in the deflection of light emitted from the astrophysical objects to the Earth due to the following two distinct effects:
@@
  • GWs or DPDM with purely gravitational interactions can induce metric perturbations, which can affect the trajectory of light emitted from the sources (Braginsky et al. (1990); Pyne et al. (1996); Kaiser & Jaffe (1997); Kopeikin et al. (1999); Book & Flanagan (2011)). As a result, the apparent positions of stars we observe slightly deviate from their initial positions. Such displacements depend on metric perturbations at both the Earth and stars, while the Earth term should be dominant, as we will explain later. It should also be mentioned that, unlike the deflection induced by ultralight scalar dark matter, which is suppressed due to an approximate spherical symmetry, the deflection caused by DPDM is generally not suppressed (Chen et al. (2023)).
@@
  • DPDM coupling to the baryon number ${\rm U}(1)_{B}$ or the baryon number minus lepton number ${\rm U}(1)_{B-L}$ can lead to the aberration effect, which also changes the apparent position of stars (Guo et al. (2019)).
@@
By precisely measuring the variation of stars’ apparent positions, it is hopeful of capturing the traces of GW and DPDM signals in the Gaia mission (Prusti et al. (2016)) and upcoming astrometric surveys like Roman (Wang et al. (2021, 2022); Haiman et al. (2023); Pardo et al. (2023)) and Theia (Boehm et al. (2017); Malbet et al. (2021); Garcia-Bellido et al. (2021)). Gaia observes positions, proper motions, and parallaxes of more than $10^{9}$ objects with unprecedented accuracy (Prusti et al. (2016)). The sensitive frequency of Gaia is determined by its survey lifetime $T_{\rm obs}=10~{}{\rm year}$ and observational cadence $\Delta t=24~{}{\rm day}$, typically, $10^{-9}~{}{\rm Hz}\lesssim f\lesssim 10^{-7}~{}{\rm Hz}$. Hence Gaia can provide complementary detection capabilities to PTA experiments for nHz astrophysical sources (Moore et al. (2017); Klioner (2018); Qin et al. (2019); O’Beirne & Cornish (2018); Mihaylov et al. (2018); Bini & Geralico (2018); Darling et al. (2018); Mihaylov et al. (2020); Guo et al. (2019); Qin et al. (2021); Aoyama et al. (2021); Jaraba et al. (2023); Liang et al. (2024); Çalışkan et al. (2024)).
@@
To extract the deflections to the apparent positions of stars caused by external signals, we need to subtract their intrinsic background motion stemming from the physical motion of stars with respect to the solar barycenter and the secular aberration due to the moving reference system. This intrinsic motion can induce large angular deflection far exceeding the resolution of Gaia. For the signals with frequencies $f\gtrsim 2/T_{\rm obs}\sim 6\times 10^{-9}~{}{\rm Hz}$, the background subtraction essentially does not influence the sensitivity, since the oscillation pattern of GWs or DPDM is distinct from the background. Conversely, if $f\lesssim 6\times 10^{-9}~{}{\rm Hz}$, a half oscillating cycle would not be observable in the time-series data within the operational lifetime of Gaia. In such cases, the signal is also partially subtracted, thereby reducing the sensitivity of detecting low-frequency GW and DPDM signals. Analogous to PTAs, the quadratic model is commonly used to subtract the background proper motion in the existing literature (Moore et al. (2017); Guo et al. (2019); Wang et al. (2021, 2022)). However, a more comprehensive study of the background motion of stars and subtraction schemes should be implemented.
@@
In this work, we propose a linear background subtraction scheme for the time-series data of apparent stellar positions in Gaia-like astrometric surveys. Specifically, we apply the least squares method to fit a linear function to the data, modeling the background noise. This fitted function is then subtracted from the data, effectively isolating the signal of interest. Upon analyzing the background motion of stars, we find that the impact on the apparent stellar position due to the constant acceleration from the galactic gravitational potential is negligibly small relative to Gaia’s sensitivity. Meanwhile, the galactocentric acceleration of the solar barycenter mimics a linear term over time in the proper motion due to aberration. These observations serve as the primary motivation for our linear subtraction approach. Nevertheless, unresolvable binary systems could bring additional noise, which cannot be effectively subtracted. We provide concrete examples by calculating the astrometric deflection induced by DPDM with purely gravitational interactions while also revisiting scenarios involving DPDM with ${\rm U}(1)_{B}$ or ${\rm U}(1)_{B-L}$ gauge symmetry and GWs from binary systems. Subsequent numerical simulation using mock data reveals that our linear subtraction scheme significantly enhances Gaia’s detection capabilities in the low-frequency range, improving sensitivity by more than an order of magnitude compared to the quadratic subtraction.
@@
The layout of this paper is as follows. In Sec. 2, we illustrate the background motion of stars can be well described by a linear model over time. In Sec. 3, we calculate astrometric deflections induced by GWs and DPDM. The numerical simulation and results are presented in Sec. 4. We summarize our main conclusions in Sec. 5.
2 Background motion of the apparent stellar locations
@@
Astrometric objects exhibit background motion around the galactic center. Their motion is randomly distributed with a potentially non-zero expectation value depending on the star’s sky location, due to the common motion of the galactic disk. We assume that the velocity $\boldsymbol{v}=(v_{r},v_{\theta},v_{\phi})$ in spherical polar coordinates relative to the galactic center follows a Gaussian distribution
@@
where $\boldsymbol{n}$ denotes the sky location of stars, $v_{0,i}(\boldsymbol{n})$ represents the expectation value of $\boldsymbol{v}$ in each direction, and $\sigma_{v,i}$ refers to the velocity dispersion. The anisotropy of the velocity dispersion has been observed by SDSS and Gaia. Approximately we have the relation $\sigma_{v,r}\simeq 2\sigma_{v,\theta/\phi}$ for stars with high metallicity, which may originate from a major merger event of a satellite galaxy between $8$ and $11$ Gyr ago (Belokurov et al. (2018)). For qualitative discussions, we assume the uncertainty $\sigma_{v}\simeq 100\,{\rm km}/{\rm s}\simeq 1.01\times 10^{-4}\,{\rm pc/year}$ for all stars on both directions on the sky sphere. At a typical distance of 1 kpc, [^1] The peak distribution of distances between the stars and us observed by Gaia is $1\sim 2$ kpc (Vallenari et al. (2023)). the angular deflection observed over $10$ years reaches $\simeq 1.01\times 10^{-6}\,{\rm rad}$, which significantly exceeds the single-exposure resolution $\sigma_{\rm ss}\simeq 100\,\mu{\rm as}=4.85\times 10^{-10}\,{\rm rad}$ of Gaia for a single star. It then becomes necessary to subtract the background proper motion from the data before searching for the GW or DPDM.
We now turn our attention to the accelerations affecting stellar and the Earth’s motion. Let us consider stars in two categories: isolated stars and the ones in binary or higher-order systems. For isolated stars, their velocities, in general, change over time. Hence, we should also estimate their accelerations, which mainly include contributions from two aspects:
  • The constant accelerations of stars. The density gradient in the Milky Way could induce constant accelerations for stars. Unlike their proper motion, the constant accelerations of stars are predictable due to the galactic mass profile, indicating the uncertainty $\sigma_{a}$ should be much smaller than the center value of the star’s acceleration, defined as $\boldsymbol{a}_{0}$. Typical stars in the Milky Way experience an acceleration of $|\boldsymbol{a}_{0}|\sim 10^{-10}\,{\rm m}/{\rm s}^{2}$ (Silverwood & Easther (2019)), which leads to a displacement of $|\boldsymbol{a}_{0}|\Delta t^{2}/2\simeq 1.6\times 10^{-10}$ pc over the 10-year observation. For a star located 1 kpc away from us, the corresponding angular deflection is about $0.03\,\mu{\rm as}$, significantly below the detection threshold.
  • The constant acceleration of solar system barycenter. Like other stellar systems, the velocity of our solar system is not precisely constant. Instead, it follows a curved orbit in the Galaxy. For the observer, the nonzero acceleration of the solar system barycenter gives rise to the aberration effect, which can be considered as extra (spurious) proper motion superimposed on stars (Kopeikin & Makarov (2006); Klioner et al. (2021))
@@
    where $\delta\boldsymbol{n}$ denotes the displacement in the apparent stellar position due to aberration, and $\boldsymbol{a}_{\rm SSB}$ is the acceleration of the solar system barycenter (SSB). The estimated value of ${\bm{a}}_{\rm SSB}$ is $2.32\times 10^{-10}~{}{\rm m}/{\rm s}^{2}$, corresponding to a proper motion amplitude of $5.05~{}\mu{\rm as}/{\rm year}$ (Klioner et al. (2021)). From Eq. (2), we realize that the displacement $\delta\boldsymbol{n}$ is linear in time, as the aberration effect should be proportional to the velocity of the observer. Therefore, although the proper motion caused by ${\bm{a}}_{\rm SSB}$ is larger than Gaia’s resolution, it can be subtracted by a function that is linearly dependent on time.
In short, the background motion of the apparent stellar locations is governed by the proper motion of stars, which can be significantly greater than the resolution of Gaia. The acceleration of stars is very minor and thus plays a negligible role. Moreover, the acceleration of the solar system barycenter yields spurious proper motion, which can be subtracted by a linear function in time. The above backgrounds can be eliminated through a linear subtraction scheme.
Next, we investigate the significant background motion from unresolved binary or higher-order systems in the Galaxy. In the practical numerical calculation, we employ a data compression technique akin to the one described in Moore et al. (2017); Klioner (2018), since the vast number of objects ($\sim 10^{9}$) in the full dataset of Gaia severely slows down the process of numerical simulation. Specifically, we compress the data of $10^{5}$ stars in the same sky area into one virtual star. For each star, the probability of being part of a binary or high-order system and having specific orbital parameters is statistically independent of the status of the other stars. After data compression, the astrometric data of $10^{5}$ stars in each sky area are added together. The background motion of individual stars in binary or higher-order systems manifests as a time-dependent noise of the compressed virtual star.
A survey of nearby ($<250$ pc) stellar objects finds $\sim 55\%$ solar-type stars are isolated, while the remaining $\sim 45\%$ stars are in binary or higher-order systems (Raghavan et al. (2010)). Here we consider the binary system consisting of two objects with masses $m_{1}$ and $m_{2}$ for simplicity. Following the Kepler’s law, the orbital period $P_{b}$ has the following relation with the semi-major axis $a_{b}$
@@
where $G$ is the gravitational constant and $q=m_{2}/m_{1}<1$. If both stars are observed by Gaia, the net contribution from the binary to the displacement of stars will approximate to
$r_{\rm net}\simeq a_{b}(1+e_{b})\frac{1-q}{1+q}\;,$ (4)
where $e_{b}$ is the orbit eccentricity. Then the net contribution will be fully canceled if $q=1$.
Figure 1: The average power spectral densities of the injected white noise (WN), proper motion (PM), and stellar binary motion, along with the associated error bars for each bin, obtained from $10^{4}$ simulations. For clarity, the error bars have been magnified by a factor of 3 to improve visibility. Green and pink dashed lines correspond to the cases with WN removed. The vertical dashed line refers to $f=2/T_{\rm obs}$ with $T_{\rm obs}=10~{}{\rm year}$ being Gaia’s observation period. The input Gaia operational parameters align with the discussion in Sec. 4.
For the $q<1$ case, we illustrate by considering a $e_{b}=0$ binary composed of two stars with the total mass $M=m_{1}+m_{2}$. The net amplitude of the observed angular deflection in one direction is
$\begin{split}A_{b}=\frac{r_{\rm net}\lambda(\iota_{b},\psi_{b})}{\sqrt{2}D_{b}}&=6.10\times 10^{-9}{\rm rad}\left(\frac{1-q}{1+q}\right)\frac{\lambda(\iota_{b},\psi_{b})}{\sqrt{2}}\\ &\times\left[\frac{P_{b}}{\rm year}\right]^{2/3}\left[\frac{D_{b}}{1{\rm kpc}}\right]^{-1}\left[\frac{M}{2M_{\odot}}\right]^{1/3}\;,\end{split}$ (5)
where $D_{b}$ is the distance between the observer and the binary barycenter and $M_{\odot}$ denotes the solar mass. Also, $\lambda(\iota_{b},\psi_{b})=\cos\iota_{b}\sin\psi_{b}+\cos\psi_{b}$ with $\iota_{b}$ and $\psi_{b}$ being the inclination angle and the polarization angle of the stellar binary plane, respectively. Eq. (5) indicates binaries with $P_{b}\gtrsim 1\,{\rm year}$ and $D_{b}\sim 1\,{\rm kpc}$ may be resolved by Gaia. According to Raghavan et al. (2010); Simonetti et al. (2020), the period distribution approximately follows the log-Normal distribution
@@
We calculate the power spectral density (PSD) originating from binary stars $S_{\rm bin}(f)$ by implementing a Monte-Carlo simulation, where we assume $45\%$ of stars are in binaries. We use Eq. (5) with $D_{b}=1$ kpc, while the orbital period is drawn from Eq. (6). Furthermore, $\cos\iota_{b}$, $\psi_{b}$ and $q$ are treated as free parameters, randomly selected from uniform distributions within the intervals of $[-1,1)$, $[0,2\pi)$ and $[0.1,1)$, respectively. In each simulation, we generate astrometric data of $10^{5}$ stars, which are then compressed into one virtual star. We calculate the PSD of astrometric data of the virtual star. The final $S_{\rm bin}(f)$ is obtained by averaging the PSD over all simulations. The total noise PSD $S_{n}(f)$ is the combination of $S_{\rm bin}(f)$ and that from the instrumental noise $S_{w}(f)$, namely,